OpenStudy (loser66):

Check my stuff, please. \[\int_1^3 2\pi x((x^2-4x+7)-4(x-2)^2)dx\]

OpenStudy (loser66):

\[2\pi \int_1^3 (-3x^3+12x^2 -3x)dx\\=6\pi \int_1^3 (-x^3+4x^2-3x) dx\] \[=6\pi*(\dfrac{-x^4}{4}+\dfrac{4x^3}{3}-\dfrac{3x^2}{2})~~from 1 ~~to~~3\]

OpenStudy (loser66):

plug the numbers in \[6\pi *[(\dfrac{-81}{4}+\dfrac{108}{3}-\dfrac{27}{2})-(\dfrac{-1}{4}+\dfrac{4}{3}-\dfrac{3}{2})]\]

OpenStudy (loser66):

I am supposed to get 16pi but I got 48/3 pi

OpenStudy (loser66):

I did \(\dfrac{-81}{4}+\dfrac{1}{4} = \dfrac{-80}{4}=-20\) \(\dfrac{108}{3}-\dfrac{4}{3}= \dfrac{104}{3}\) \(dfrac{-27}{2}+\dfrac{3}{2}=\dfrac{-24}{2}=-12\)

OpenStudy (loser66):

Add them all I got a wrong answer :(

OpenStudy (unklerhaukus):

\[\int_1^32\pi x\big[(x^2-4x+7)-4(x-2)^2\big]\,\mathrm dx\\ =2\pi\int_1^3x\big[x^2-4x+7-4(x^2-4x+4)\big]\,\mathrm dx\\ =2\pi\int_1^3x\big[x^2-4x+7-4x^2+16x-16\big]\,\mathrm dx\\ =2\pi\int_1^3 x\big[-3x^2+12x-9\big]\,\mathrm dx\\ =2\pi\int_1^3 [-3x^3+12x^2-9x]\,\mathrm dx\\ =\dots\]

OpenStudy (loser66):

Yes, Sir

OpenStudy (alekos):

Looks like Wolf ram is wrong

OpenStudy (loser66):

No, not that!! I let 2pi out , just put the integral to calculate, it gives me 8, then *2pi =16pi that is the answer from the back of the book

OpenStudy (unklerhaukus):

\[=2\pi\int_1^3 [-3x^3+12x^2-9x]\,\mathrm dx\\ =6\pi\int_1^3 [-x^3+4x^2-3x]\,\mathrm dx\\ =6\pi\left.\Big[-\frac14x^4+\frac43x^3-\frac32x^2\Big]\right|_1^3\\ =\dots\]

OpenStudy (loser66):

Yes

OpenStudy (unklerhaukus):

\[6\pi\Big[\left(\frac{-81}{4}+\frac{108}{3}-\frac{27}{2}\right)-\left(\frac{-1}{4}+\frac{4}{3}-\frac{3}{2}\right)\Big]\\ =6\pi\Big[\big(-20.5+36-13.5\big)-\left(-0.25+1\tfrac{1}{3}-1.5\right)\Big]\\ =6\pi\Big[\big(2.25\big)-\left(-0.75+\tfrac{1}{3}\right)\Big]\\ =6\pi\Big[2.25+0.75-\tfrac{1}{3}\Big]\\ =6\pi\Big[3-1/3\Big]\\ =\pi[18-2]\\=\dots \]

OpenStudy (loser66):

oh man!!! Thank you so much. stupid me!! hihihihi

OpenStudy (unklerhaukus):

[the wolfram link forgot the 2π]

OpenStudy (unklerhaukus):

all the brackets negative signs and fractions make errors likely

OpenStudy (loser66):

I didn't put 2pi in, it's good.

OpenStudy (loser66):

Again, thank you "uncle" hihihihi

OpenStudy (unklerhaukus):

hiii!!!